# Advancing Universal Deep Learning for Electronic-Structure Hamiltonian Prediction of Materials

### v2

@@ 1 Introduction
Understanding the electronic structure is fundamental to unraveling how electrons govern the properties of condensed matter systems. This knowledge is essential for predicting a wide range of material characteristics, such as electrical conductivity, magnetism, optical behavior, and chemical activity, which are vital for technologies spanning from electronics to sustainable energy and advanced catalysis. At the heart of these calculations is the challenge of determining the system’s Hamiltonian matrix, whose eigenvalues and eigenstates yield important quantities like energy levels, band structures, and electronic wavefunctions. Traditionally, Density Functional Theory (DFT) (Hohenberg & Kohn, 1964; Kohn & Sham, 1965) has been the standard approach for these problems. However, as shown in Fig. 1 (a), DFT relies heavily on the self-consistent (SC) procedure, which demands repeated (denoted as $T$ turns), computationally intensive diagonalizations of large matrices—each scaling as $\mathcal{O}(N^{3})$ with system size $N$, making simulations of large or complex materials extremely resource-consuming. Recently, deep learning has emerged as a powerful tool in the physical sciences (Zhang et al., 2025). As shown in Fig. 1 (b), modern deep neural network methods (Gong et al., 2023; Yu et al., 2023b; Zhang et al., 2024; Wang et al., 2024b; Li et al., 2025; Yin et al., 2025) can predict Hamiltonians, i.e., the core physical quantities in electronic structure calculations, directly from atomic configurations in an efficient way, circumventing the computationally expensive SC loop and dramatically accelerating computations. This paradigm shift lowers the computational barriers associated with electronic structure calculations, unlocking the simulation and design of unprecedentedly large-scale materials systems, driving new innovation in materials discovery and engineering. Please refer to Appendix A.1 for background introduction.
@@
However, deep learning methods still face substantial challenges in achieving accurate and generalizable Hamiltonian prediction, primarily due to the extremely complex and fundamentally difficult nature of the input–output mapping that the neural network must learn, making it difficult to generalize across diverse material systems. Consequently, it has become common practice to constrain the scope, such as limiting the range of supported elements, neglecting spin–orbit coupling (SOC) effects, or reducing the number of orbitals considered, as discussed in Section 2. While such strategies help alleviate modeling burdens, they also restrict the applicability of these methods to the full diversity and complexity of real-world materials. What’s more, large open-source materials datasets for the training and evaluation of general Hamiltonian learning models are also rare.
@@
First, we dive deeply into the traditional DFT computational process outlined in Appendix A.1 and novely introduce a physical quantity that helps mitigate the complexity of the input–output mapping encountered by deep learning models for Hamiltonian prediction. This quantity is the zeroth-step Hamiltonian $\mathbf{H}^{(0)}$, which is efficiently constructed from the initial electron density $\rho^{(0)}(\mathbf{r})$, given by the sum of the charge densities of isolated atoms, without the requirement of matrix diagonalization. As $\mathbf{H}^{(0)}$ efficiently encodes essential information about the system’s electronic structure, we use it as one of the input features to the neural network. Unlike existing methods that rely on randomly initialized atom and edge embeddings lacking physical prior knowledge, $\mathbf{H}^{(0)}$ provides richer physical context by embedding the intrinsic characteristics of diverse elements into a unified representation space, thereby enabling robust generalization across chemically complex material systems. Furthermore, we make our method a correction approach by taking $\Delta\mathbf{H}=\mathbf{H}^{(T)}-\mathbf{H}^{(0)}$ instead of $\mathbf{H}^{(T)}$ as the regression target of neural networks, significantly reducing their output space. This compression of the output space helps the regression model to finely fit and generalize.
@@
Second, we present a network architecture that strictly adheres to E(3) symmetry while maintaining high non-linear expressiveness for Hamiltonian prediction by extending the TraceGrad (Yin et al., 2025) method to Transformer framework, thereby providing ample capacity for flexible and accurate modeling of atomic systems for Hamiltonian prediction across a wide range of elements in the periodic table. Furthermore, we introduce model ensemble techniques to enhance the capacity of the framework for handling complex scenarios in Hamiltonian prediction.
Third, we propose a joint optimization framework that simultaneously refines both real-space (R-space) and reciprocal-space (k-space) Hamiltonians. Most existing methods regress only the real-space Hamiltonian, but the large condition number of the overlap matrix can amplify errors in predicted eigenvalues and eigenfunctions, leading to suboptimal physical fidelity. By jointly optimizing R- and k-space Hamiltonians with a unified loss, our approach ensures accurate prediction of key electronic properties, e.g. the band structures.
@@
On the dataset side, we curate a diverse-collection large benchmark dataset, Materials-HAM-SOC, containing $17,000$ material structures generated using DFT softwares. The dataset spans $68$ elements from the first six rows of the periodic table and explicitly incorporates spin–orbit coupling (SOC) effects. To ensure the accuracy of the DFT calculations, we employ high-quality pseudopotentials that include as many valence electrons as possible, enabling our model to handle physically complex and highly challenging systems. We adopt high-quality atomic orbital basis sets, up to 4s2p2d1f orbitals for each element, to ensure fine-grained description of electronic structures. This dataset establishes a challenging yet comprehensive benchmark for evaluating generalization across chemically and structurally diverse systems.
@@
2 Related Work
The foundation of deep learning-based electronic-structure Hamiltonian prediction lies in constructing neural networks that respect E(3)-symmetry, including invariance to translations, as well as invariance and equivariance to rotations (elements of SO(3)) and inversions (Geiger & Smidt, 2022). Since traditional non-linear activation functions, when applied directly to SO(3)-equivariant features, may break equivariance, a central research topic is to reconcile strong non-linear expressiveness with strict SO(3)-equivariance. An early attempt to address this problem was the use of gated activation functions (Weiler et al., 2018), which first apply non-linear activations to SO(3)-invariant features and then use them as coefficients to scale the SO(3)-equivariant features. Representative works adopting this mechanism include DeepH-E3 (Gong et al., 2023) and QHNet (Yu et al., 2023b) for Hamiltonian prediction. To further enhance non-linear expressiveness of E(3)-networks, Zitnick et al. (2022) and Passaro & Zitnick (2023) proposed eSCN (equivariant Spherical Channel Networks), which applies non-linear operations to the coefficients obtained from the spherical decomposition of features. This approach has been widely used in Hamiltonian prediction (Wang et al., 2024a; b; Yu et al., 2025) tasks. Nevertheless, eSCN methods project features onto discrete basis functions via inner-product operations, degrading strict SO(3)-equivariance to equivariance with respect to only a discrete sub-group. Moreover, they replace SO(3) convolutions with SO(2) convolutions, sacrificing strict inversion equivariance. Such compromises may limit the physical soundness of the results. Yin et al. (2025) proposed the TraceGrad method, which introduces an SO(3)-invariant supervision signal, i.e., the trace quantity, constructed from Hamiltonian labels, to supervise high-fidelity non-linear SO(3)-invariant features, and propagates non-linearity to SO(3)-equivariant features via a gradient-based mechanism derived from the SO(3)-invariant ones. This method effectively unifies strict E(3)-symmetry, with strong non-linear expressiveness for Hamiltonian prediction; however, the backbone network it adopted is a simple graph neural network and has not yet evolved into a non-linear equivariant Transformer framework.
@@ 2 Related Work
Despite these progresses, deep learning methods for Hamiltonian prediction still face substantial challenges on generalization performance, which can be summarized as follows. First, crystalline materials commonly found in nature can be composed of over $65$ different elements from the first six rows of the periodic table, leading to an exceptionally large and heterogeneous input space for deep neural network models. Existing deep learning methods for Hamiltonian prediction typically employ learnable embeddings to represent nodes (atoms) and edges (atom pairs). These embeddings are randomly initialized and learned directly from the dataset, without incorporating any explicit physical priors. As a result, they struggle to capture the fundamental physical relationships between different atoms and across different material systems, which are crucial for generalization. Second, as illustrated in Figure 1, the regression target, namely the self-consistent electronic-structure Hamiltonian, is inherently high-dimensional and complex, especially when considering SOC effects. For instance, a system containing several tens of atoms may involve nearly several thousands of non-zero Hamiltonian matrix elements that need to be accurately predicted. Most of the existing methods attempt to directly predict the entire self-consistent Hamiltonian matrix, namely $\mathbf{H}^{(T)}$ as formulated in Appendix A.1, placing a heavy burden on the model due to the vast size of the output space, often resulting in optimization difficulties during training and limited generalization to unseen systems. In addition, most existing methods treat the real-space Hamiltonian as the sole regression target, which can lead to sub-optimal physical fidelity in down-stream applications, particularly in capturing low-energy band structures accurately. Although Li et al. (2025) designed a method for molecular systems to reduce the regression space of Hamiltonians and introduced a basis transformation of the Hamiltonian matrix on the loss function to improve the prediction accuracy of down-stream physical quantities, their approach is limited to molecular systems and is not applicable to periodic crystalline materials.
As a result, constructing a unified model that generalizes across diverse crystal prototypes remains challenging, and many existing approaches explicitly constrain their scope. For example, Li et al. (2022), Gong et al. (2023), and Xia et al. (2025) each train and evaluate their methods within a single material system (e.g., MoS2, Bi2Se3, or a-HfO2), without assessing cross-material generalization. More recently, DeepH-2 (Wang et al., 2024b) broadened coverage to systems involving elements primarily from the first four rows of the periodic table; however, they neglected SOC and reduced the orbital basis by omitting $f$-orbitals. While such choices help reduce computational and modeling complexity, they may limit broad applicability to the full diversity of real-world materials. Zhong et al. (2025) developed a Hamiltonian prediction model aimed at a broader range of element types, while also highlighting the challenge of achieving consistently high accuracy across diverse crystal systems. Moreover, open-source datasets with a broad and diverse collection of materials dedicated to training and validating universal Hamiltonian models across the periodic table remain scarce. Although the QH9 (Yu et al., 2023a) dataset is a well-known open-source Hamiltonian dataset, it consists of molecular systems rather than periodic material systems, and includes only structures composed of C, H, O, N, and F elements. To solve these challenges, this work presents an advanced unified deep learning framework together with a large benchmark dataset for Hamiltonian prediction, targeting broader generalization across richer classes of materials.
@@ 3 Method
3.1 Input Descriptors and Output Targets
As shown at the lower part of Fig. 2, we use the displacement vector-based descriptors between atoms that lie within the cutoff distance, together with the zeroth-step Hamiltonian (calculated as detailed in Appendix A.1), as the input features for the neural network. Introducing the zeroth-step Hamiltonian as the input features for the neural network is one of the core innovations of our framework. The zeroth-step Hamiltonian $\mathbf{H}^{(0)}$, derived from the initial charge density $\rho^{(0)}(\mathbf{r})$, obtained as the sum of neutral atomic charges, reflects the information of different elements in the system, including the strength of the electron-ion interactions (pseudopotential) and a preliminary estimate of the electron-electron interactions. These components directly influence the system’s electronic structure. This method encodes the characteristics of different elements into a unified representation space, which enhances the model’s ability to generalize across diverse material systems.
The zeroth-step Hamiltonian $\mathbf{H}^{(0)}$ can be decomposed into on-site sub-matrices, which represent the Hamiltonian blocks corresponding to each atom and its own orbitals, and off-site sub-matrices, which capture the interactions between different atoms. These two types of sub-matrices can naturally serve as the initial descriptors for nodes and edges, respectively, in the graph neural network.
Moreover, as detailed in Appendix A.1, computing the zeroth-step Hamiltonian requires no matrix diagonalization, so its cost scales with the number of non-zero matrix elements: it is approximately $\mathcal{O}(N^{2})$ for small systems with $N$ atoms and asymptotically approaches $\mathcal{O}(N)$ for sufficiently large systems as the neighbor count saturates for each atom. This matches the scaling behavior of message passing mechanism of graph neural networks, ensuring that incorporating the zeroth-step Hamiltonian as a new input descriptor does not worsen the $\mathcal{O}$-asymptotics.
On the output side of the neural network, our approach predicts the correction term $\Delta\mathbf{H}=\mathbf{H}^{(T)}-\mathbf{H}^{(0)}$. This residual formulation reframes the learning task as a refinement problem, where the network adjusts the physically informed zeroth-step Hamiltonian $\mathbf{H}^{(0)}$ toward its self-consistent counterpart. By significantly reducing both the dimensionality and numerical range of the regression target, this approach simplifies the learning problem and enables the model to focus on capturing only the essential differences rather than reconstructing the entire Hamiltonian from scratch. As a result, it promotes generalization performance across diverse atomic configurations.
@@ 3.2 Neural Network Architecture
We present a Transformer architecture that not only maintains strict E(3)-symmetry but also achieves strong non-linear expressiveness, as shown in Fig. 2. Our E(3)-symmetry graph attention mechanism is developed from Equiformer (Liao & Smidt, 2023). While Equiformer was designed for regression tasks where the target quantity is essentially a node-level atomic property (e.g., force fields), our Hamiltonian target is fundamentally an edge-level property defined on atomic pairs. This distinction necessitates stronger modeling of edge features and motivates the development of our attention mechanism. First, we explicitly maintain and update edge features across multiple layers, rather than generating them only temporarily from node features on demand (Liao & Smidt, 2023). In this way, the computation of attention weights incorporates both the node features and the persistently maintained edge features. Second, motivated by the decay behavior of Hamiltonian matrix elements with respect to interatomic distance, we explicitly incorporate interatomic distances by introducing distance embeddings as additional signals in the computation of attention weights, enabling the model to better exploit distance information for inference. Third, the attention weights between nodes are directly applied to update edge features via multiplicative operations, and are subsequently refined through equivariant transformations. Together, these developments substantially enhance the capacity of the model to represent edge features, from which the Hamiltonian is regressed.
@@
As analyzed in Section 2, the TraceGrad method (Yin et al., 2025) can maintain strong non-linear expressiveness while preserving strict E(3) symmetry. We extends TraceGrad into the Transformer framework for electronic-structure Hamiltonian prediction. As shown in the middle of Fig. 2, for an atomic pair $(a,b)$, the SO(3)-equivariant edge feature $\mathbf{f}^{{}^{\prime}\text{(edge)}}_{ab}$ output by the E(3)-symmetry multi-head graph attention module is fed into the TraceGrad module to produce the non-linear SO(3)-invariant feature $z^{(\text{edge})}_{ab}$, which is subsequently passed to the SO(3)-invariant decoder and trained under the supervision of the SO(3)-invariant trace quantity $\mathbf{T}=\mathrm{tr}(\Delta\mathbf{H}\cdot\Delta\mathbf{H}^{\dagger})$. The learned non-linear expressiveness in $z^{(\text{edge})}_{ab}$ is subsequently delivered into the equivariant feature by $\mathbf{o}^{\text{(edge)}}_{ab}=\mathbf{f}^{{}^{\prime}\text{(edge)}}_{ab}+\frac{\partial\,z^{(\text{edge})}_{ab}}{\partial\,\mathbf{f}^{{}^{\prime}\text{(edge)}}_{ab}}$, where $\mathbf{o}^{\text{(edge)}}_{ab}$ represents the non-linearity-enhanced SO(3)-equivariant edge feature, which, together with the node feature, are fed into the subsequent encoding modules of the Transformer as well as the SO(3)-equivariant decoder to regress the correction term $\Delta\mathbf{H}$.

### v3

@@ 1 Introduction
Understanding the electronic structure is fundamental to unraveling how electrons govern the properties of condensed matter systems. This knowledge is essential for predicting a wide range of material characteristics, such as electrical conductivity, magnetism, optical behavior, and chemical activity, which are vital for technologies spanning from electronics to sustainable energy and advanced catalysis. At the heart of these calculations is the challenge of determining the system’s Hamiltonian matrix, whose eigenvalues and eigenstates yield important quantities like energy levels, band structures, and electronic wavefunctions. Traditionally, Density Functional Theory (DFT) (Hohenberg and Kohn, 1964; Kohn and Sham, 1965) has been the standard approach for these problems. However, as shown in Fig. 1 (a), DFT relies heavily on the self-consistent (SC) procedure, which demands repeated (denoted as $T$ turns), computationally intensive diagonalizations of large matrices, each scaling as $\mathcal{O}(N^{3})$ with system size $N$, making simulations of large or complex materials extremely resource-consuming. Recently, deep learning has emerged as a powerful tool in the physical sciences (Zhang et al., 2025). As shown in Fig. 1 (b), modern deep neural network methods (Gong et al., 2023; Yu et al., 2023b; Zhang et al., 2024; Wang et al., 2024c; Li et al., 2025; Yin et al., 2025) can predict Hamiltonians, i.e., the core physical quantities in electronic structure calculations, directly from atomic configurations in an efficient way, circumventing the computationally expensive SC loop and dramatically accelerating computations. This paradigm shift lowers the computational barriers associated with electronic structure calculations, unlocking the simulation and design of unprecedentedly large-scale materials systems, driving new innovation in materials discovery and engineering. Please refer to Appendix A for background introduction.
@@
However, deep learning methods still face substantial challenges in achieving accurate and generalizable Hamiltonian prediction, primarily due to the extremely complex and fundamentally difficult nature of the input–output mapping that the neural network must learn, making it difficult to generalize across diverse material systems. Consequently, it has become common practice to constrain the scope, such as limiting the range of supported elements, neglecting spin–orbit coupling (SOC) effects, or reducing the number of orbitals considered, as thoroughly discussed in Appendix C. While such strategies help alleviate modeling burdens, they also restrict the applicability of these methods to the full diversity and complexity of real-world materials. What’s more, large open-source materials datasets for the training and evaluation of general Hamiltonian learning models are also rare.
@@
First, we dive deeply into the traditional DFT computational process outlined in Appendix A and introduce a physical quantity that helps mitigate the complexity of the input–output mapping encountered by deep learning models for Hamiltonian prediction. This quantity is the zeroth-step Hamiltonian $\mathbf{H}^{(0)}$, which is efficiently constructed from the initial electron density $\rho^{(0)}(\mathbf{r})$, given by the sum of the charge densities of isolated atoms, without the requirement of matrix diagonalization. As $\mathbf{H}^{(0)}$ efficiently encodes essential information about the system’s electronic structure, we innovatively incorporate it as one of the input features to the neural network. Unlike existing methods that rely on randomly initialized atom and edge embeddings, which lack physical prior knowledge and may suffer from issues of sparsity as analyzed in Appendix M, $\mathbf{H}^{(0)}$ provides richer physical context by embedding the intrinsic characteristics of diverse elements into a unified representation space, thereby enabling robust generalization across chemically complex material systems. Moreover, inspired by the delta-learning paradigm (Bowman et al., 2022), we predict the correction term $\Delta\mathbf{H}=\mathbf{H}^{(T)}-\mathbf{H}^{(0)}$ rather than directly predicting the entire $\mathbf{H}^{(T)}$, reducing both the dimensionality and numerical range of the regression target, and enabling the model to focus on capturing only the essential differences rather than reconstructing the entire Hamiltonian from scratch.
@@
Second, we present a network architecture that strictly adheres to E(3)-symmetry while maintaining high non-linear expressiveness for Hamiltonian prediction by extending the TraceGrad (Yin et al., 2025) method to Transformer framework, thereby providing ample capacity for flexible and accurate modeling of atomic systems for Hamiltonian prediction across a wide range of elements in the periodic table. Furthermore, we introduce model ensemble techniques to enhance the capacity of the framework for handling complex scenarios in Hamiltonian prediction.
Third, we propose a joint optimization framework that simultaneously refines both real-space (R-space) and reciprocal-space (k-space) Hamiltonians. Most existing methods regress only the real-space Hamiltonian, but the large condition number of the overlap matrix can amplify errors in predicted eigenvalues and eigenfunctions, leading to suboptimal physical fidelity. Although recent work (Li et al., 2025) has explored strategies to mitigate this error amplification, their attempts are tailored to finite molecular systems, and also overlook the inherent gauge freedom in Hamiltonian representations (Wang et al., 2024c). In contrast, our method explicitly targets the decoupling of energy subspaces in k-space for infinite periodic systems to eliminate spurious “ghost states” and strictly preserve the band topology. Furthermore, we resolve the gauge ambiguity by analytically determining the optimal gauge parameter within our joint optimization framework, thereby stabilizing the optimization landscape and ensuring unique, physically consistent predictions.
@@
On the dataset side, we curate a diverse-collection large benchmark dataset, Materials-HAM-SOC, containing $17,000$ material structures generated using DFT softwares. The dataset spans more than $60$ elements from the first six rows of the periodic table and explicitly incorporates spin–orbit coupling (SOC) effects. To ensure the accuracy of the DFT calculations, we employ high-quality pseudopotentials that include as many valence electrons as possible, enabling our model to handle physically complex and highly challenging systems. We adopt high-quality atomic orbital basis sets, up to 4s2p2d1f orbitals for each element, to ensure fine-grained description of electronic structures. This dataset establishes a challenging yet comprehensive benchmark for evaluating generalization across chemically and structurally diverse systems.
@@ 2 Method
2.1 Input Descriptors
As shown at the lower part of Fig. 2, we use the displacement vector-based descriptors between atoms that lie within the cutoff distance, together with the zeroth-step Hamiltonian (calculated as detailed in Appendix A), as the input features for the neural network. Introducing the zeroth-step Hamiltonian as the input features for the neural network is one of the core innovations of our framework. The zeroth-step Hamiltonian $\mathbf{H}^{(0)}$, derived from the initial charge density $\rho^{(0)}(\mathbf{r})$, obtained as the sum of neutral atomic charges, reflects the information of different elements in the system, including the strength of the electron-ion interactions (pseudopotential) and a preliminary estimate of the electron-electron interactions. These components directly influence the system’s electronic structure. By embedding $\mathbf{H}^{(0)}$ as the inputs, our method encodes the characteristics of different elements into a unified representation space, bringing a powerful capability to generalize across diverse material systems.
The zeroth-step Hamiltonian $\mathbf{H}^{(0)}$ can be decomposed into on-site sub-matrices, which represent the Hamiltonian blocks corresponding to each atom and its own orbitals, and off-site sub-matrices, which capture the interactions between different atoms. These two types of sub-matrices can naturally serve as the initial descriptors for nodes and edges, respectively, in the graph neural network, as detailed in Appendix E.
What’s more, as detailed in Appendix A, computing the zeroth-step Hamiltonian requires no matrix diagonalization, so its cost scales with the number of non-zero matrix elements: it is approximately $\mathcal{O}(N^{2})$ for small systems with $N$ atoms and asymptotically approaches $\mathcal{O}(N)$ for sufficiently large systems as the neighbor count saturates for each atom. This matches the scaling behavior of the message passing mechanism of graph neural networks, ensuring that incorporating the zeroth-step Hamiltonian as a new input descriptor does not worsen the $\mathcal{O}$-asymptotics.
@@ 2.2 Neural Network Architecture
Accurate Hamiltonian prediction requires the neural network to strictly adhere to the symmetries of the E(3) group. While translation symmetry can be easily implemented using relative coordinates, maintaining O(3)-equivariance while also achieving significant expressive power presents a challenging and fundamental problem. To provide the necessary background on the directly relevant basic concepts, we refer the reader to Appendix B-D.
@@
We present a Transformer architecture that not only maintains strict E(3)-symmetry but also achieves strong non-linear expressiveness, as shown in Fig. 2. Our E(3)-symmetry graph attention mechanism is developed from Equiformer (Liao and Smidt, 2023). While Equiformer was designed for regression tasks where the target quantity is essentially a node-level atomic property (e.g., force fields), our Hamiltonian target is fundamentally an edge-level property defined on atomic pairs. This distinction necessitates stronger modeling of edge features and motivates the development of our attention mechanism. First, we explicitly maintain and update edge features across multiple layers, rather than generating them only temporarily from node features on demand (Liao and Smidt, 2023). In this way, the computation of attention weights incorporates both the node features and the persistently maintained edge features. Second, motivated by the decay behavior of Hamiltonian matrix elements with respect to interatomic distance, we explicitly incorporate interatomic distances by introducing distance embeddings as additional signals in the computation of attention weights, enabling the model to better exploit distance information for inference. Third, the attention weights between nodes are directly applied to update edge features via multiplicative operations, and are subsequently refined through equivariant transformations. Together, these developments substantially enhance the capacity of the model to represent edge features, from which the Hamiltonian is regressed.
@@
As analyzed in Section C, the TraceGrad mechanism (Yin et al., 2025) can maintain strong non-linear expressiveness while preserving strict E(3)-symmetry  [^1] Although the original paper of TraceGrad emphasizes its SO(3)-equivariance, it is straightforward to prove that it also preserves E(3)-symmetry, including translation-invariance and O(3)-equivariance in this context.. We extend TraceGrad into the Transformer framework for electronic-structure Hamiltonian prediction. As shown in the middle of Fig. 2, for an atomic pair $(a,b)$, the O(3)-equivariant edge feature $\mathbf{f}^{{}^{\prime}\text{(edge)}}_{ab}$ is fed into the TraceGrad module to produce the non-linear O(3)-invariant feature $z^{(\text{edge})}_{ab}$, which is subsequently passed to the O(3)-invariant decoder and trained under the supervision of the O(3)-invariant trace quantity $\mathbf{T}=\mathrm{tr}(\Delta\mathbf{H}\cdot\Delta\mathbf{H}^{\dagger})$. The learned non-linear expressiveness in $z^{(\text{edge})}_{ab}$ is subsequently delivered into the equivariant feature by $\mathbf{o}^{\text{(edge)}}_{ab}=\mathbf{f}^{{}^{\prime}\text{(edge)}}_{ab}+\frac{\partial\,z^{(\text{edge})}_{ab}}{\partial\,\mathbf{f}^{{}^{\prime}\text{(edge)}}_{ab}}$, where $\mathbf{o}^{\text{(edge)}}_{ab}$ represents the non-linearity-enhanced O(3)-equivariant edge feature, which, together with the node feature, are fed into the subsequent encoding modules of the Transformer followed by the $O(3)$-equivariant decoder and a Wigner–Eckart converter (Gong et al., 2023) to regress the correction term $\Delta\mathbf{H}=\mathbf{H}^{(T)}-\mathbf{H}^{(0)}$. This residual formulation defines the learning task as a delta-learning problem, reducing both the dimensionality and numerical range of the regression target.